# Rethinking Approximate Gaussian Inference in Classification

### v1

@@ Abstract
In classification tasks, softmax functions are ubiquitously used as output activations to produce predictive probabilities. Such outputs only capture aleatoric uncertainty. To capture epistemic uncertainty, approximate Gaussian inference methods have been proposed, which output Gaussian distributions over the logit space. Predictives are then obtained as the expectations of the Gaussian distributions pushed forward through the softmax. However, such softmax Gaussian integrals cannot be solved analytically, and Monte Carlo (MC) approximations can be costly and noisy. We propose a simple change in the learning objective which allows the exact computation of predictives and enjoys improved training dynamics, with no runtime or memory overhead. This framework is compatible with a family of output activation functions that includes the softmax, as well as element-wise normCDF and sigmoid. Moreover, it allows for approximating the Gaussian pushforwards with Dirichlet distributions by analytic moment matching. We evaluate our approach combined with several approximate Gaussian inference methods (Laplace, HET, SNGP) on large- and small-scale datasets (ImageNet, CIFAR-10), demonstrating improved uncertainty quantification capabilities compared to softmax MC sampling. Code is available at github.com/bmucsanyi/probit.
Deep learning, Uncertainty quantification, Probabilistic modeling, Bayesian inference
[…]
[ linewidth=1pt, roundcorner=3pt, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=-5pt, innerbottommargin=4pt, leftmargin=0pt, rightmargin=0pt, skipabove=2pt, skipbelow=0pt, ]eqbox
@@ 1 Introduction
Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set. This synthetic dataset consists of 100 logit means $\bm{\mu}$ and standard deviations $\bm{\sigma}$, whose components are i.i.d. uniformly distributed. For a fair comparison we scale the logit dataset for the different activations to match the activations to first order, so that $\mu_{c}\in[-1,1]$, $\sigma_{c}\in[0,1]$ for sigmoid, $\mu_{c}\in[-1/2-\log 2,1/2-\log 2]$, $\sigma_{c}\in[0,1/2]$ for softmax, $\mu_{c}\in[-\sqrt{\pi/8},\sqrt{\pi/8}]$, $\sigma_{c}\in[0,\pi/8]$ for normCDF. The MC approximations are capped at a fixed computational budget of $10000$ class samples, i.e. $\lceil 10000/C\rceil$ samples. We see that, even without enforcing the constraint Equation 18, our approximations are the only ones that do not become worse with the number of classes. See Appendix C for theoretical analyses of the quality of the different approximations.
@@
With the usual assumption of i.i.d. data, the likelihood under the model $\bm{h}$ is given by ($\delta$ is the Kronecker delta)
$p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N})$ $=\prod_{n=1}^{N}p(c_{n}\mid x_{n})$ (3)
$=\prod_{n=1}^{N}\prod_{c=1}^{C}h_{c}(x_{n})^{\delta_{c_{n},c}}.$
This yields a natural loss function for classification problems, the cross-entropy (CE) loss
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\log(p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N}))$ (4)
$=-\sum_{n=1}^{N}\sum_{c=1}^{C}\delta_{c_{n},c}\log(h_{c}(x_{n})).$
For typical models $\bm{h}$ trained through such approximate maximum likelihood estimation, an output probability vector $\bm{h}(x)\in\Delta^{C-1}$ should be interpreted as an estimate of $p_{\text{gen}}(\bm{y}\mid x)$, the probability under the generative model. Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data. In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty (Hüllermeier & Waegeman, 2021).
The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (as in “a distribution over probability distributions”). That is, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ (Sensoy et al., 2018; Charpentier et al., 2020).
@@
where $f$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax. For example, Heteroscedastic Classifiers (HET) (Collier et al., 2021) learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network. In linearised Laplace approximations (Daxberger et al., 2021), only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters. Last-layer Laplace approximations (Kristiadi et al., 2020) are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network. Spectral-Normalized Gaussian processes (SNGP) (Liu et al., 2020) are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
To obtain predictive probabilities from a distributional classifier as in Equation 5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\int_{\Delta^{C-1}}p\;\mathrm{h}(x)(dp)$ (6)
$=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(y)\;\mathrm{f}(x)(dy).$
In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts (Depeweg et al., 2018; Mukhoti et al., 2023; Wimmer et al., 2023). These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$. Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula (Atchison & Shen, 1980), this does not seem of practical use. Even predictives cannot be computed analytically. On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes. Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited. This limits the application of such Gaussian inference methods in classification tasks with large number of classes, such as those common in computer vision and natural language processing.
@@ 1.1 Proposed Recipe
We can summarise our framework as a 4-step recipe:
  1. 1.
    Choose an approximate Gaussian inference method: a method that outputs logit covariances.
  2. 2.
    Choose an output activation for which the corresponding Gaussian integral can be solved (Section 2).
  3. 3.
    Train with an objective that enforces a certain normalisation constraint (Section 4).
  4. 4.
    At inference time, use our analytic formulas (see Section 2 for the predictives, Section 3 for other quantities of interest).
2 Gaussian Inference with Analytic Predictives
@@ 2.1 Motivation from Softmax Models
A key difficulty in obtaining exact or closed-form approximate predictive probabilities of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ is that pushforwards of Gaussian distributions through the softmax are intractable. The softmax function is the composition $\bm{\operatorname{\mathbf{softmax}}}=\bm{n}\circ\bm{\exp}$, where $\bm{n}$ is the normalisation function given by $\bm{n}(\bm{q}):=\bm{q}/(q_{1}+\dots+q_{C})$ and $\bm{\exp}$ is applied element-wise. Thus, $\mathrm{h}$ can be written as the composition (c.f. Equation 5)
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\exp}_{*}}\mathcal{P}(\mathbb{R}_{>0}^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (7)
Now it is easy to see that (Section A.1)
@@
$\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]$ $=\int_{\mathbb{R}^{C}}\bm{\exp}(\bm{y})\;\mathrm{f}(x)(d\bm{y})$ (9)
$=\bm{\exp}\left(\bm{\mu}(x)+\frac{\bm{\sigma}^{2}(x)}{2}\right)$
where $\bm{\mu}(x)$ and $\bm{\sigma}^{2}(x)$ are, respectively, the mean and the diagonal of the covariance of $\mathrm{f}(x)$, and the vector division is element-wise. Note that no approximations are made so far.
We can use this to approximate the predictive of the model in closed form:
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]$ (10)
$\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]}{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]}$
$=\frac{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\exp_{*}\mathrm{f}_{c}(x)}\left[Q_{c}\right]}$
@@ 2.2 The General Framework
In fact, we can apply the previous argumentation to a general family of output activation functions: suppose we can write the output activation as $\bm{n}\circ\bm{\varphi}$, where $\bm{\varphi}$ is the element-wise application of some activation function $\varphi\colon\mathbb{R}\to\mathbb{R}_{>0}$. So we have (c.f. Equation 7)
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\varphi}_{*}}\mathcal{P}(\mathbb{R}_{>0}^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (11)
If we can solve the one-dimensional integrals (c.f. Equation 8)
$\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (12)
then we can compute the approximate predictive (c.f. Equation 10)
@@
For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section A.2),
$\int_{\mathbb{R}}\Phi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)$ (14)
@@
Taking instead $\varphi=\rho$, the logistic sigmoid, we can approximate Equation 12 with the probit approximation (Section A.3)
@@
2.3 Making the Approximations Exact
At this point, the approximate Equation 13 may seem ad-hoc. However, note that Equation 13 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable, as can be seen from the derivation in Equation 10. Thus, we would like to enforce a constraint in the model of the form
$\sum_{c=1}^{C}Q_{c}$ $=s(x)$ $\text{where }\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)$ (18)
where $s\colon\mathcal{X}\to\mathbb{R}$ is a deterministic function. In other words, the law of $\sum_{c=1}^{C}Q_{c}$ is a Dirac measure for all $x$.
In neural networks, constraints can be learned through appropriate choices of learning objectives during training, ensuring that they are satisfied within the training data distribution. In Section 4, we describe how this can be achieved depending on $\bm{\varphi}$ and the method that is used to obtain the mapping to Gaussian distributions $\mathrm{f}$.
Given the constraint of Equation 18 is satisfied, Equations 10 and 15 ($\varphi=\exp\text{ or }\Phi$) both become exact. For Equation 17 ($\varphi=\rho$), the only error incurred is the one from the probit approximation (Equation 16) per each class. In the synthetic experiment (Figure 1), we see that this induces no significant effect on the quality of the predictive. We justify this observation with the following result:

### v2

@@ Abstract
In classification tasks, softmax functions are ubiquitously used as output activations to produce predictive probabilities. Such outputs only capture aleatoric uncertainty. To capture epistemic uncertainty, approximate Gaussian inference methods have been proposed. We develop a common formalism to describe such methods, which we view as outputting Gaussian distributions over the logit space. Predictives are then obtained as the expectations of the Gaussian distributions pushed forward through the softmax. However, such softmax Gaussian integrals cannot be solved analytically, and Monte Carlo (MC) approximations can be costly and noisy. We propose to replace the softmax activation by element-wise normCDF or sigmoid, which allows for the accurate sampling-free approximation of predictives. This also enables the approximation of the Gaussian pushforwards by Dirichlet distributions with moment matching. This approach entirely eliminates the runtime and memory overhead associated with MC sampling. We evaluate it combined with several approximate Gaussian inference methods (Laplace, HET, SNGP) on large- and small-scale datasets (ImageNet, CIFAR-100, CIFAR-10), demonstrating improved uncertainty quantification capabilities compared to softmax MC sampling.
@@ 1 Introduction
With the usual assumption of i.i.d. data, the likelihood under the model $\bm{h}$ is given by
$p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N})$ $=\prod_{n=1}^{N}p(c_{n}\mid x_{n})=\prod_{n=1}^{N}\prod_{c=1}^{C}h_{c}(x_{n})^{\delta_{c_{n},c}},$ (3)
where $\delta$ is the Kronecker delta. This yields a natural loss function for classification problems, the cross-entropy (CE) loss
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\log(p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N}))=-\sum_{n=1}^{N}\sum_{c=1}^{C}\delta_{c_{n},c}\log(h_{c}(x_{n})).$ (4)
For typical models $\bm{h}$ trained through such approximate maximum likelihood estimation, an output probability vector $\bm{h}(x)\in\Delta^{C-1}$ should be interpreted as an estimate of $p_{\text{gen}}(\bm{y}\mid x)$, the probability under the generative model. Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data. In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [15].
Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set. This synthetic dataset consists of 100 logit means $\bm{\mu}$ and standard deviations $\bm{\sigma}$, whose components are i.i.d. uniformly distributed. For a fair comparison we scale the logit dataset for the different activations to match the activations to first order, so that $\mu_{c}\in[-1,1]$, $\sigma_{c}\in[0,1]$ for sigmoid, $\mu_{c}\in[-1/2-\log 2,1/2-\log 2]$, $\sigma_{c}\in[0,1/2]$ for softmax, $\mu_{c}\in[-\sqrt{\pi/8},\sqrt{\pi/8}]$, $\sigma_{c}\in[0,\pi/8]$ for normCDF. The MC approximations are capped at a fixed computational budget of $10000$ class samples, i.e. $\lceil 10000/C\rceil$ samples (see also Section C.1).
The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (that is ‘a distribution over probability distributions’). Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [32, 4].
@@
where $\mathrm{f}$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax. For example, Heteroscedastic Classifiers (HET) [5] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network. In linearised Laplace approximations [8], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters. Last-layer Laplace approximations [16] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network. Spectral-Normalized Gaussian processes (SNGP) [20] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
To obtain predictive probabilities from a distributional classifier as in Eq. 5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\int_{\Delta^{C-1}}p\;\mathrm{h}(x)(dp)=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(y)\;\mathrm{f}(x)(dy).$ (6)
In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [10, 28, 39]. These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$. Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [1], this does not seem of practical use—even predictives cannot be computed analytically. On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes. Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited. This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modeling.
@@ 1.1 Proposed Recipe
We summarise our framework as a 3-step recipe:
  1. (i)
    Choose an approximate Gaussian inference method that outputs logit means and covariances.
  2. (ii)
    Choose an output activation that yields a closed-form Gaussian integral (Section 2).
  3. (iii)
    At inference time, use our closed-form formulas (see Section 2 for the predictives, Section 3 for other quantities of interest).
2 Gaussian Inference with Closed-Form Predictives
@@ 2.1 Motivation from Softmax Models
A key difficulty in obtaining exact or closed-form approximate predictive probabilities of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ is that pushforwards of Gaussian distributions through the softmax are intractable. The softmax function is the composition $\bm{\operatorname{\mathbf{softmax}}}=\bm{n}\circ\bm{\exp}$, where $\bm{n}$ is the normalisation function given by $\bm{n}(\bm{q}):=\bm{q}/(q_{1}+\dots+q_{C})$ and $\bm{\exp}$ is applied element-wise. Thus, $\mathrm{h}$ can be written as the composition (see Eq. 5)
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\exp}_{*}}\mathcal{P}((0,\infty)^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (7)
Now it is straightforward to see that (Section A.1)
@@
$\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]$ $=\int_{\mathbb{R}^{C}}\bm{\exp}(\bm{y})\;\mathrm{f}(x)(d\bm{y})=\bm{\exp}\left(\bm{\mu}(x)+\frac{\bm{\sigma}^{2}(x)}{2}\right)$ (9)
where $\bm{\mu}(x)$ and $\bm{\sigma}^{2}(x)$ are, respectively, the mean and the diagonal of the covariance of $\mathrm{f}(x)$, and the vector division is element-wise. Note that no approximations have been made so far.
One can use this to approximate the model’s predictive in closed form:
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]=\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]$ $\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]}{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]}$ (10)
@@ 2.2 The General Framework
The previous argumentation can be applied to a general family of output activation functions: suppose the output activation is of the form $\bm{n}\circ\bm{\varphi}$, where $\bm{\varphi}$ is the element-wise application of some activation function $\varphi\colon\mathbb{R}\to(0,\infty)$. Then we have (see Eq. 7)
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\varphi}_{*}}\mathcal{P}((0,\infty)^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (11)
Solving the one-dimensional integrals $\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (see Eq. 8) allows us to obtain the approximate predictive (see Eq. 10)
@@
For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section A.2)
$\int_{\mathbb{R}}\Phi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right).$ (13)
@@
Taking instead $\varphi=\rho$, the logistic sigmoid, we approximate Eq. 13 using the probit approximation (Section A.3),
@@
2.3 Theoretical Analysis of the Approximation’s Quality
In Fig. 1, we see empirically that the quality of our predictive approximations outperforms existing predictive approximations on synthetic data. We now explore the quality of our predictive approximations theoretically. We conduct the analysis for the case $\varphi=\rho$. Let
$q=q(\mu,\sigma^{2}):=\mathbb{E}_{Q\sim\rho_{*}\mathcal{N}(\mu,\sigma^{2})}[Q]\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}(\mu,\sigma^{2})=\hat{q},$ (17)
where $\hat{q}$ is the probit approximation (Section A.3). Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$ where $(\mu_{c},\sigma_{c}^{2})$ is the logit mean-variance pair for class $c$, $\bm{p}$ for the true predictive and $\bm{\hat{p}}$ for our approximate predictive, i.e.
$\bm{p}:=\mathbb{E}_{\bm{P}\sim(\bm{n}\circ\bm{\rho})_{*}\mathcal{N}(\bm{\mu},\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2}))}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}.$ (18)
To analyse the quality of the approximation, we thus want to bound $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$. Clearly, when all the $Q_{c}$ are deterministic (i.e. constant random variables), where $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)$, we have $\bm{p}=\bm{\hat{p}}$. Alternatively, from the derivation Eq. 10, we see that in the limiting case Eq. 12 is exact when $\sum_{c=1}^{C}Q_{c}$ is deterministic. This is a weaker assumption than requiring all $Q_{c}$ to be deterministic individually. In fact, we can derive a rate of decay of $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$ as a function of $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)$: